# Uncovering Autoregressive LLM Knowledge of Thematic Fit in Event Representation

### v1

@@ Abstract
[…]
@@
Safeyah Khaled Alshemali Columbia University New York, NY, USA ska2145@columbia.edu                        Daniel Bauer Columbia University New York, NY, USA db2711@columbia.edu                        Yuval Marton University of Washington WA, USA ymarton@uw.edu
@@ 1 Introduction
Do large language models (LLMs) have consistent, expressible knowledge (including linguistic knowledge) about the thematic fit of a given event argument to its assigned semantic role for a given predicate (the event ‘trigger’)? A naive approach of prompting LLMs might yield negative results and thus lead to wrong conclusions about the absence of such knowledge. However, we hypothesize that this knowledge does exist, at least to a certain degree, but needs eliciting. We test our hypothesis along three axes: (1) Whether structured reasoning in prompting improves the model’s estimations, following the success of chain-of-thought  prompting (Wei et al., 2022). (2) Whether providing generated sentences as contexts helps, compared to providing the model with only the lemma of the predicate, the lemma of the argument’s syntactic head, and its semantic role. (3) Whether asking for categorical output (such as “low” or “high” thematic fit) yields better results than asking for a numeric output $(0..1)$ as the model’s thematic fit estimation.
@@
The thematic fit estimation task measures the compatibility between the predicate (typically a verb), its argument (typically a noun phrase), and a specific semantic role assigned to the argument. [^1] Thematic fit is related to, but separate from, semantic role labeling (SRL). Given a sentence, SRL should output the semantic frames in the sentence, including the predicates, arguments, and their semantic roles in the frame. Thematic fit takes the predicates, arguments, and roles as input, and outputs a score. Understanding the thematic fit of individual arguments/roles is thought to be required for SRL, but is not usually modeled explicitly in SRL systems. For example, the verb ‘eat’ invokes a set of potential arguments for specific semantic roles. If the role is Agent (usually denoted as Arg0  in PropBank; see sec.§3.2), then ‘people’, ‘husband’, etc., would be a good fit, as opposed to ‘pizza’, ‘apple’, etc., as the latter group would better fit a Patient role (often denoted as Arg1  in PropBank). If the role is Location, then ‘restaurant’, ‘kitchen’, etc., would be a good fit due to their properties. If the role is Instrument, then ‘fork’, ‘knife’, etc. would be a good fit due to their use as tools while eating. And so on. Psycholinguistic experiments show that humans rely on generalized, prototypical knowledge about events and their typical participants during language understanding, early in the process and prior to syntactic interpretation (McRae and Matsuki, 2009; Bicknell et al., 2010). The thematic fit task examines how well computational event representations are aligned with this type of human event knowledge.
@@
The thematic fit task is challenging due to the lack of direct training data. Thematic fit norms (see §3.2), which summarize human judgments, provide a benchmark for evaluation, but are not large enough for direct supervised training. Instead, thematic fit models are traditionally trained on proxy data (either PropBank (Kingsbury and Palmer, 2002) or FrameNet (Johnson et al., 2016) annotations). Recently, pretrained language models have shown state-of-the-art performance without task-specific supervised training on many computational semantics tasks, such as synonym judgment (Levy et al., 2017), similarity judgments (Hill et al., 2015), and more. However, on thematic fit, these models, including masked LMs such as BERT (Devlin et al., 2019), have so far only yielded modest improvements over previous approaches (Pedinotti et al., 2021). In this work, we examine if more recent, large autoregressive models, such as GPT (Radford et al., 2018; Achiam et al., 2023), can produce better results using prompting. We attempt to elicit their relevant linguistic knowledge along the following three axes:
Step-by-Step Prompting
@@
Several studies show LLMs perform remarkably well in challenging NLP tasks such as Planning (Wang et al., 2022), automatic scoring (Lee et al., 2024), and question answering (Wang et al., 2023), when using ‘chain-of-thought  prompting’ (Wei et al., 2022). This technique breaks a problem into sub-steps toward answering the core question. It seems to help LLMs in a similar way a scratchpad would help someone solve a math exercise. We view thematic fit estimation as a linguistic reasoning task because it can be broken into clear sub-steps: listing the argument’s (salient or relevant) properties, listing the (prototypical or necessary) properties required for the role “assigned” by the given predicate, and estimating the argument’s thematic fit to the assigned role given these two lists. Thus, we compare the LLMs’ performance using a “naive” prompting approach (Simple Prompting hereafter) with the chain-of-thought  approach (Step-by-Step Prompting  hereafter) on the thematic fit estimation task.
Generated Sentences vs. Lemma Tuples
The test sets we use in this study only provide tuples of lemmas, for example, (predicate: eat, argument: pizza, role: Arg0/Agent), and a thematic fit score (an average of several human raters’ scores). Since LLMs were trained on full sentences (as opposed to lemma tuples), we assume this format will be unnatural for them, even if the relevant linguistic and/or world knowledge exists. Therefore, we attempt to reconstruct full sentences with these predicates, arguments, and roles. Furthermore, upon a close reading, we find that the human raters were indeed provided full sentences for the rating (McRae et al., 1998), but this fact seems to have been lost over the years, and only the lemma tuple format has been used (Tilk et al., 2016; Hong et al., 2018; Muthupari et al., 2022). We close here a full circle, bringing the conditions closer to the original.
Predefined Categories Vs. Numeric Score
@@
The thematic fit ratings in our test sets are Likert scale ratings (following McRae et al. (1998)) averaged over 7 human raters. Thus, a straightforward evaluation would prompt the model to output numeric scores as well. For each test set, we then calculate the Spearman Rank Correlation Coefficient (§3.4) LLMs  were designed to handle textual data, so outputting numeric scores may lead to inconsistent results. In fact the LLMs’ numeric scores are actually a textual output, which we then convert to its corresponding numeric value. We, therefore, also prompt the LLMs to score the thematic fit according to predefined categories, which we then convert to corresponding predefined values as well (and then compute a correlation score as before).
@@
  • First study to use prompting autoregressive LLMs to directly estimate thematic fit.
  • Comparison of Simple Prompting  vs. Step-by-Step Prompting  for thematic fit estimation.
  • Comparison of generated sentences as additional input context vs. only using lemma forms of the predicate and argument’s syntactic head.
  • Comparison of numeric outputs vs. predefined categories (“Low”, “High”, etc.)
  • New state-of-the-art with our GPT-powered method on all tested tasks (see details of fair comparison to the baselines in Sect. §4).
@@ 2 Related Work
Early distributional approaches to the thematic fit task compute the similarity between vector representations of an argument (usually for the argument’s syntactic head word, obtained from a distributional semantic model or static word embeddings) and a prototype vector for the predicate/role pair (Erk et al., 2010; Baroni and Lenci, 2010; Santus et al., 2017; Sayeed et al., 2015; Greenberg et al., 2015). Our first baseline (hereafter BG) is the distributional approach by Greenberg et al., which is based on the syntactically structured distributional semantic model by Baroni and Lenci. More recent approaches use neural networks to learn event representations. Tilk et al. (2016) describe a non-incremental role filler model (NN RF, baseline B0) trained to predict role fillers, given as input a predicate, target role, and optionally other contextual roles and their fillers. Hong et al. (2018) improve over Tilk et al. by adding role prediction as a multi-task objective, and improving pooling of context roles/fillers (ResRoFA-MT, baseline B1).Marton and Sayeed (2021) replicate Hong et al. (2018)’s model and train it on higher quality silver labels, experimenting with the quality/quantity tradeoff of annotations (baseline B2). Muthupari et al. (2022) focus on evaluating the capability of various models, or model elements, to capture thematic fit. They compare initializing the Hong et al. model with random vectors vs. pre-trained GloVe embeddings (Pennington et al., 2014), and other modifications. We add three of their higher performing models to our baselines: (B3a) Glove-shared-not-tuned, (B3b) Glove-shared-tuned, (B3c) Glove-shared-tuned-20%M. B3a and B3b report the average scores over several runs, whereas B3c is identical to B3b except for reporting the maximum score over several runs, all trained on the same 20% subset of the data.
Contemporary work using pretrained LLMs (without any additional supervised training) for thematic fit either examines if these models can differentiate plausible from implausible sentences (’the teacher bought the laptop’ vs. ’the laptop bought the teacher’) (Kauf et al., 2023), or used BERT-style models to predict role fillers for a masked argument in the context of a generated sentence Metheniti et al. (2020); Pedinotti et al. (2021). Notably, Pedinotti et al. only report a moderate improvement over a distributional approach that uses a high quality structured distributional model (Chersoni et al., 2019). Finally, Testa et al. (2023) examine the interaction between thematic fit and the ability of LLMs to resolve verbal ellipsis. We are the first to use Simple and Step-by-Step Prompting with autoregressive LLMs directly for thematic fit.
@@ 3 Experiments
The experiments were designed to investigate three axes with two possible values each. Therefore, we conduct $82^{3}=$ 8 experiments (see Table 1), labeled according to these axes. Each experiment has a prompting template according to the input, output, and reasoning settings, as detailed in Appendix §A.
Reasoning Form Input Form Output Form
Simple Prompting Head Lemma Tuples 1.1 Num. Score
1.2 Cat. Score
Generated Sentences 2.1 Num. Score
2.2 Cat. Score
Step-by-Step Prompting Head Lemma Tuples 3.1 Num. Score
3.2 Cat. Score
Generated Sentences 4.1 Num. Score
4.2 Cat. Score
Figure 1: Simple Prompting Approach
Figure 2: Step-by-Step Prompting Approach
3.1 Experiment Design: Three Axes
Axis1: Reasoning Form: Simple Prompting  vs. Step-by-Step Prompting
@@ 3.1 Experiment Design: Three Axes
Axis 2: Input Form: Head Lemma Tuples vs. Generated Sentences
@@
Axis 3: Output Form: Categorical vs. Numeric Thematic Fit Scoring
Each experiment may generate one of two forms of output: a numeric thematic fit score (Exp.x.1) or a predefined thematic fit category (Exp.x.2). The numeric thematic fit score (numeric score, hereafter) is a numeric value in the range [0,1]. The predefined thematic fit categories (categorical score, hereafter) are ’Near-Impossible’, ’Low’, ’Medium’, ’High’, or ’Near-Perfect’. These output score categories are then converted to the following numeric values: (0.0, 0.25, 0.5, 0.75, 1.0), respectively, in a post-processing script, in order to compute a correlation score compared to the human ratings.
@@ 3.2 Datasets and Evaluation
Training: We used pre-trained LLMs (see sec. §3.3) “off the shelf” without modification or additional training. In an ideal world, we would be able to point the reader to the full pre-training data. Test sets: We used the following datasets to evaluate our methods.
McRae et al. (1998), hereafter ’McRae’, contains $1,444$ predicate-argument-role-rating tuples, developed with the participation of 120 native English speakers undergraduate students at the University of Rochester. Students were asked to rate the role typicality (thematic fit) of a predicate (verb) and the argument (noun) in sentence fragments on a Likert scale of 1 to 7. The public dataset does not contain the sentence fragments. The roles are PropBank’s Arg0  and Arg1 (Kingsbury and Palmer, 2002). We removed eight duplicate records.
Padó et al. (2006), hereafter ’Pado’, contains 414 tuples in the same form as McRae et al. (1998); however, the role could also be Arg2. To prevent bias in rating the same verb-noun with different roles, one hundred volunteer English native speakers judged these tuples in four separate randomly ordered lists.
Ferretti et al. (2001) evaluate thematic fit for the roles Instrument  (hereafter, the Fer-Ins dataset) and Location  (hereafter, Fer-Loc dataset). These datasets were done with the input from fifty-eight students. The rating scale is the same as in McRae et al. (1998). However, for Instrument, the participants were asked, “How common is it for someone to use each of the following to perform the eating action?” For Location, the question was, “How common is it for someone to eat in each of the following locations?”. We removed the indefinite articles attached to the argument in the Fer-Loc dataset, in order to consider only the lemma of the syntactic head of the argument (not the phrase).
Evaluation: We used Spearman’s Rank Correlation (Spearman, 1904) between each of the test sets (each set sorted by the average human ratings for each item) and the corresponding model output (sorted as well). Categorical output was converted to numeric values (detailed in §3.1).

### v2

@@ Abstract
We show closed models possess much thematic fit knowledge and set a new state of the art, while open models also seem to capture much relevant knowledge (in semantic filtering), but yield lower scores. Surprisingly, multi-step reasoning only helped closed models (with few exceptions); generated sentences hurt closed models’ performance; and output form had little to no effect. We analyze the reasons for these findings, and conclude that more foundational work is needed for a single LLM to perform the best on all tasks with the same experimental condition, let alone improve results further. Source code is available at: https://github.com/SafeyahShemali/LLM_Thematic_Fit_25
@@
Safeyah Khaled Alshemali Imperial College London London, UK ska24@ic.ac.uk          Daniel Bauer Columbia University New York, NY, USA db2711@columbia.edu          Yuval Marton University of Washington WA, USA ymarton@uw.edu
@@ 1 Introduction
Do large language models (LLMs) have consistent, expressible knowledge (including linguistic knowledge) of event arguments’ thematic fit to their assigned semantic role for a given predicate? A naive approach of prompting LLMs might yield negative or weak results. However, we hypothesize that this knowledge does exist, at least to a certain degree, but needs eliciting. We test our hypothesis along three axes: (1) Reasoning: Does structured reasoning in prompting improve the model’s estimations? [^1] Following chain-of-thought (wei2022chain). (2) Input: Does providing generated sentences as contexts help, compared to providing only a triplet of keywords? [^2] Traditionally consisting of the predicate’s lemma, the argument’s syntactic head’s lemma, and its semantic role. (3) Output: Does specifying categorical output (such as “low” or “high”) yield better results than a numerical output $(0..1)$ as the model’s thematic fit estimation?
@@
The thematic fit estimation task measures the compatibility between the predicate (typically a verb), its argument (typically a noun phrase), and a specific semantic role assigned to the argument. [^3] Thematic fit is related to, but separate from, semantic role labeling (SRL). Given a sentence, SRL should output the semantic frames in the sentence, including the predicates, arguments, and their semantic roles in the frame. Thematic fit takes the predicates, arguments, and roles as input, and outputs a score. Understanding the thematic fit of individual arguments/roles is thought to be required for SRL, but is not usually modeled explicitly in SRL systems. For theoretical and historical details, see gruber1965studies; fillmore1968lexical; dowty1991thematic; parsons:1990; mcrae1998modeling inter alia. For example, the verb ‘eat’ invokes a set of potential arguments for specific semantic roles. If the role is Agent (usually denoted as Arg0 in PropBank; see sec.§3.2), then ‘people’, ‘cat’, etc., would be a good fit, as opposed to ‘pizza’, ‘apple’, etc., as the latter group would better fit a Patient role (often denoted as Arg1 in PropBank). If the role is Location, then ‘restaurant’, ‘kitchen’, etc., would be a good fit due to their properties. If the role is Instrument, then ‘fork’, ‘knife’, etc. would be a good fit due to their use as tools while eating. And so on. Psycholinguistic experiments show that humans rely on generalized, prototypical knowledge about events and their typical participants during language understanding, early in the process and prior to syntactic interpretation (mcrae2009people; bicknell2010effects).
@@
Thematic fit estimation applies data-driven event representations and computational methods to mimic, or align with this type of human knowledge of events. This is challenging due to the lack of direct labeled training data. Thematic fit norms (see §3.2), which summarize human judgments (aka ratings), provide a benchmark for evaluation, but are not large enough for direct supervised training. Instead, thematic fit models are traditionally trained on proxy data (either PropBank (kingsbury2002treebank) or FrameNet (johnson2016framenet) annotations). Recently, pretrained language models have shown state-of-the-art performance without task-specific supervised training on many computational semantics tasks, such as synonym judgment (levy2017semantic), similarity judgments (hill2015simlex), and more. However, on thematic fit, these models, including masked LMs such as BERT (devlin2019bert), have so far only yielded modest improvements over previous approaches (pedinotti2021cat). In this work, we examine if more recent, large autoregressive models, such as GPT (classicalGPT; radford2018improving; achiam2023gpt), can produce better results via prompting. We attempt to elicit their relevant linguistic knowledge along the following three axes:
Axis 1: Reasoning (Simple vs. Step-by-Step Prompting)
@@
Several studies show LLMs perform remarkably well in challenging NLP tasks such as Planning (wang2022towards), automatic scoring (lee2024applying), and question answering (wang2023keqing), when using ‘chain-of-thought prompting’ (wei2022chain). This technique breaks a problem into sub-steps toward answering the core question. It seems to help LLMs in a similar way a scratchpad would help someone solve a math exercise. [^4] It proved so influential that several leading labs have developed ‘thinking’ models with ’test-time compute’, incorporating this approach in the post-training and settings. We view thematic fit estimation as a linguistic reasoning task because it can be broken into clear sub-steps (see §3.1). Thus, we compare the LLMs’ performance using a “naive” single-prompt approach (Simple Prompting hereafter) with the chain-of-thought approach (Step-by-Step Prompting hereafter) on the thematic fit estimation task.
Axis 2: Input (Generated Sentences vs. Lemma Tuples)
[…]
Axis 3: Output (Predefined Categories vs. Numeric Score)
@@
The thematic fit ratings in our test sets are Likert scale ratings (following mcrae1998modeling) averaged over 7 human raters. Thus, a straightforward evaluation would prompt the model to output numeric scores as well. However, LLMs were designed to handle textual data, so outputting numeric scores may lead to inconsistent results. Moreover, the LLMs’ numeric scores are actually a textual output, which we then convert to its corresponding numeric value. Therefore, we also prompt the LLMs to rate the thematic fit according to predefined categories (“Low”, “High”, etc.), which we then convert to corresponding predefined numeric values for evaluation purposes (see Section §3.2). [^6] We leave for future work whether it is useful to use LLMs’ logprobs for this task, and how to best do that.
@@
  • First study to prompt autoregressive LLMs in order to directly estimate thematic fit.
  • New state-of-the-art thematic fit estimation.
  • Comparison of three axes: reasoning form, input form, and output form.
  • A novel semantic filtering process to exclude semantically incoherent or semantically incompatible generated sentences.
  • Showing and analyzing the weakness of new vs. older GPT models in this task, and why chain-of-thought is not always helpful here.
@@ 2 Related Work
Early distributional approaches to the thematic fit task compute the similarity between vector representations of an argument (usually for the argument’s syntactic head word, obtained from a distributional semantic model or static word embeddings) and a prototype vector for the predicate/role pair (erk2010flexible; baroni2010distributional; santus2017; sayeed2015exploration; greenberg2015improving). Our first baseline (hereafter BG) is the distributional approach by greenberg2015improving, which is based on the syntactically structured distributional semantic model by baroni2010distributional. More recent approaches use neural networks to learn event representations. tilk2016event describe a non-incremental role filler model (NN RF, baseline B0) trained to predict role fillers, given as input a predicate, target role, and optionally other contextual roles and their fillers. hong2018learning improve over tilk2016event by adding role prediction as a multi-task objective, and improving pooling of context roles/fillers (ResRoFA-MT, baseline B1). Marton2021ThematicFB replicate hong2018learning’s model and train it on higher quality silver labels, experimenting with the quality/quantity tradeoff of annotations (baseline B2). muthupari2022s focus on evaluating the capability of various models, or model elements, to capture thematic fit. They compare initializing the hong2018learning model with random vectors vs. pre-trained GloVe embeddings (pennington2014glove), and other modifications. We add three of their higher performing models to our baselines: (B3a) Glove-shared-not-tuned, (B3b) Glove-shared-tuned, (B3c) Glove-shared-tuned-20%M. B3a and B3b report the average scores over several runs, whereas B3c is identical to B3b except for reporting the maximum score over several runs, all trained on the same 20% subset of the data.
Contemporary work using pretrained LLMs (without any additional supervised training) for thematic fit either examines if these models can differentiate plausible from implausible sentences (’the teacher bought the laptop’ vs. ’the laptop bought the teacher’) (kauf2023event), or used BERT-style models to predict role fillers for a masked argument in the context of a generated sentence metheniti2020relevant; pedinotti2021cat. Notably, pedinotti2021cat only report a moderate improvement over a distributional approach that uses a high quality structured distributional model (chersoni2019structured). Finally, testa2023we examine the interaction between thematic fit and the ability of LLMs to resolve verbal ellipsis.
We are the first to use prompting with autoregressive LLMs directly for thematic fit.
Figure 1: Experiment Method. For more details of the Step-by-Step Prompting and Semantic Filtering, see §3.1. The output of Exp.3.x - Exp.4.x contains the score’s justification in addition to the score.
Reasoning Form Input Form Output Form Exp
Simple Prompting Head Lemma Tuples Num. Score 1.1
Cat. Score 1.2
Generated Sentences Num. Score 2.1
Cat. Score 2.2
Step-by-Step Prompting Head Lemma Tuples Num. Score 3.1
Cat. Score 3.2
Generated Sentences Num. Score 4.1
Cat. Score 4.2
@@ 3 Experiments
The experiments were designed to investigate three axes (§1) with two possible values each. Therefore, we conduct $2^{3}=8$ experiments. We refer to them as ’Exp $n.m$’ (see Table 1). Each experiment has a prompting template according to the reasoning, input, and output settings, as detailed in Appendix §A. Consistent with the order in the introduction, we will next detail the design for the reasoning form, the input form, and the output form.
3.1 Method: Reasoning, Input and Output
Axis 1: Reasoning (Simple Prompting vs. Step-by-Step Prompting)
@@ 3.1 Method: Reasoning, Input and Output
Axis 2: Input (Lemma Tuples vs. Generated Sentences)
@@
The sentence must pass all the steps of the filtering process in at least one trial, or it will be filtered out. For each of the five sentences, the final fit score is averaged. If steps (1) or (2) fail, we back off to using the lemma tuples. The second column in Table 1 specifies the form of the input.
Axis 3: Output (Categorical vs. Numeric Scoring)
Each experiment may generate one of two forms of output: a numeric thematic fit score (Exp.x.1 in Table 1) or a predefined thematic fit category (Exp.x.2). The numeric thematic fit score (numeric score, hereafter) is a numeric value in the range [0,1]. The predefined thematic fit categories (categorical score, hereafter) are ’Unlikely/Impossible’, ’Low’, ’Medium’, ’High’, or ’Perfect’. These output score categories are then converted to the following numeric values: (0.0, 0.25, 0.5, 0.75, 1.0), respectively, in a post-processing script, for evaluation purposes (see §3.2). In Exp.3.x and Exp.4.x, the model is asked to generate a justification for the score it assigns to the input, based on its previous Step-by-Step Prompting reasoning output.
@@ 3.2 Datasets and Evaluation
Training set:
We used pre-trained LLMs (see sec. §3.4) “off the shelf” without modification. In an ideal world, we would be able to point the reader to the full pre-training data. We did not add psycholinguistic (or other) labeled data to the training set, nor did we fine-tune the LLMs.
Test sets:
[…]
mcrae1998modeling, hereafter ’McRae’, contains $1,444$ predicate-argument-role-rating tuples, developed with the participation of 120 native English speakers undergraduate students at the University of Rochester. Students were asked to rate the role typicality (thematic fit) of a predicate (verb) and the argument (noun) in sentence fragments on a Likert scale of 1 to 7. The public dataset does not contain the sentence fragments. The roles are PropBank’s Arg0 and Arg1 (kingsbury2002treebank). We removed eight duplicate records.
pado2006modelling, hereafter ’Pado’, contains 414 tuples in the same form as mcrae1998modeling; however, the role could also be Arg2. To prevent bias in rating the same verb-noun with different roles, one hundred native English speaker volunteers judged tuples in four randomly ordered lists.
ferretti2001integrating evaluate thematic fit for the roles Instrument (hereafter, the Fer-Ins dataset) and Location (hereafter, Fer-Loc dataset). These datasets were done with the input from fifty-eight students. The rating scale is the same as in mcrae1998modeling. However, for Instrument, the participants were asked, “How common is it for someone to use each of the following to perform the eating action?” For Location, the question was, “How common is it for someone to eat in each of the following locations?”. We removed the indefinite articles attached to the argument in the Fer-Loc dataset, in order to consider only the lemma of the syntactic head of the argument (not the phrase).
Evaluation:
For each test set, we used Spearman’s Rank Correlation Coefficient ($\rho$) (spearman:1904) between two lists: the test set, sorted by the average human ratings for each item, and the corresponding model output (sorted as well). Categorical output was converted to numeric values, so it can be sorted similarly too before conducting this evaluation (see §3.1).
3.3 Hyperparametes
We conducted a preliminary study to determine good temperature and top_p parameter values for the LLMs. We used GPT-3.5 on a subset of the Fer-Ins dataset (20 samples $\sim$ 8%) and ran a parameter sweep over temperature values (0.0, 0.5, 0.9) and top_p values (0.5, 0.7, 0.95). Temperature values were chosen near-uniformly across the scale to examine a wide range of the model creativity setting. Higher values correspond to higher randomness (aka "creativity") in the output. High values of top_p limit the randomness in the output to more probable tokens. Since we were more interested in accurate output, and less interested in creative writing for our purposes here, we tested top-p threshold values on the higher half of the scale. Temperature 0.0 and top_p 0.95 showed the highest correlation coefficient with human judgments on our sample. We set the max_tokens parameter based on the length of the prompt and its expected output. For experiments with Lemma Tuples input, we set max_tokens to 100, whereas for Generated Sentences input, we set it to 300 (for Simple prompting). As Step-by-Step Prompting is expected to be lengthy to clarify the underlying reasons, we set max_tokens to 600 there. We used these settings in all experiments.